OpenStudy (anonymous):

determine dx/dy: 1)y=(sec x+3 ln tanx)^1/4 2)y=lnsqrte^3x(x^2-1)

OpenStudy (callisto):

@mkrugel Are you familiar with the chain rule?

OpenStudy (callisto):

For question one, y=(sec x+3 ln tanx)^1/4 Let u = sec x+3 ln tanx So, y = u^(1/4) \[\frac{dy}{dx} = \frac{dy}{du} \times \frac{du}{dx}\] Find dy/du and du/dx respectively. Can you do it now?